OpenStudy (anonymous):

Of the 106 first-degree murder cases from 2002 through the first half of 2004 in a certain superior court, 6 cases were thrown out of the system, 70 cases were plea-bargained, and 30 cases went to trial. What is the probability that a case selected at random has the following outcomes? (Round your answers to three decimal places.) (a) settled through plea bargaining (b) went to trial

OpenStudy (anonymous):

\[\frac{70}{106}\] i guess

OpenStudy (anonymous):

how do i find part a and b

Directrix (directrix):

Probability of an outcome = the number of favorable outcomes divided by the number of possible outcomes P(Plea Bar) = 70/106 because there were 70 outcomes of plea bargaining and 106 total outcomes. Note: favorable outcome does not mean something necessarily "good." It means an outcome favorable to the probability you seek. P(Trial) = 30/106 for the same reasons. Note: This type probability is known as empirical probability. The probabilities are not fixed but determinied by the data given.

OpenStudy (anonymous):

thanks alot!!